OpenStudy (anonymous):

What is the derivative of (x^2)/(x^3 +1) and why?

OpenStudy (amistre64):

Dx(t/b) = (bt'-b't)/b^2

OpenStudy (amistre64):

or Dx(t/b) = Dx(t b^-1) = t' b^-1 + -b^-2 t

OpenStudy (anonymous):

thank you sir!

OpenStudy (anonymous):

or ma'am

OpenStudy (amistre64):

:) yw

OpenStudy (amistre64):

2x(x^3 +1)- (x^2)3x^2 --------------------- ; simplify as needed (x^3 +1)^2

OpenStudy (angela210793):

[(x^2)/(x^3 +1)]'=(-x^4+1)/(x^3+1)^2

OpenStudy (angela210793):

(U/V)'=9U'V-UV')/V^2

OpenStudy (amistre64):

uvs are fine and all; but I perfor t and b to distiguish top and bottom easier

OpenStudy (amistre64):

like the product rule: r'l + rl' , not that its hard to keep track of those lol

OpenStudy (angela210793):

ok but i meant Dx(t/b) = (bt'-b't)/b^2 must b b't-bt'/b^2

OpenStudy (amistre64):

nope; bt' is first.... always has been you bring the bottom up to derive the top

OpenStudy (amistre64):

just like yours ;)

OpenStudy (angela210793):

let me see smth pls :)

OpenStudy (amistre64):

(U/V)'=U'V-UV')/V^2 t/b = t'b - tb' / b^2 t/b = bt' - b't / b^2

OpenStudy (amistre64):

its all the same ;)

OpenStudy (angela210793):

it is? O.o

OpenStudy (angela210793):

i mean in this case u subtract them...not like...U*V which would b ok if u switched places...

OpenStudy (amistre64):

nobody switched places; they all 3 are identical

OpenStudy (amistre64):

u'v = t'b = bt'

OpenStudy (angela210793):

oops :$ :$ ...i got it.....i thought u had written smth else...sorry ^_^

OpenStudy (amistre64):

uv' = tb' = b't

OpenStudy (angela210793):

I got it :)